# Minimax optimal differentially private synthetic data for smooth queries

### v1

@@ 1 Introduction
Much of the existing DP literature focuses on designing task-specific mechanisms tailored to a predefined workload (e.g., a fixed set of queries or a particular model class). While often highly accurate, these approaches can require substantial expertise, ranging from sensitivity analysis and mechanism selection (e.g., Laplace/Gaussian noise, exponential mechanism [26]) to careful privacy accounting across iterative computations. A complementary and increasingly practical alternative is differentially private synthetic data generation, where one releases a synthetic dataset that approximates statistical properties of the original data while satisfying differential privacy [32, 9, 61, 8]. By the post-processing property [26] of differential privacy, analysts can subsequently perform a wide range of downstream tasks on the released synthetic dataset without incurring additional privacy loss.
@@ 1.1 Differentially private synthetic data
Within this Wasserstein framework, [11] gave the first $\varepsilon$-DP algorithm achieving the minimax rate $(\varepsilon n)^{-1/d}$ (up to polylogarithmic factors) for data in the hypercube $0,1]^{d}$ under the $\ell_{\infty}$ metric. Subsequently, [37] developed two classes of algorithms that attain the same minimax-optimal $(\varepsilon n)^{-1/d}$ rate for $d\geq 2$. These constructions have enabled the analysis of several downstream learning tasks [31, 62, 18]. Under the relaxed notion of $(\varepsilon,\delta)$-DP, [48] analyzed a Chebyshev moment matching algorithm that achieves the same $n^{-1/d}$ Wasserstein rate.
Recent work has sought to surpass the worst-case $n^{-1/d}$ scaling by exploiting additional structure in the data. For example, [36] considered datasets concentrated near a $d^{\prime}$-dimensional subspace and obtained a rate $(\varepsilon n)^{-1/d^{\prime}}$ up to polynomial factors in $d$, while [23] studied data supported on an unknown $d^{\prime}$-dimensional manifold in $0,1]^{d}$. When the dataset satisfies suitable sparsity conditions, [39] showed that the error rate can be further improved via an adaptation of the Private Measure Mechanism introduced in [37]. In a related direction, [29] analyzed the practical Private Evolution algorithm [45] through its connection to the Private Signed Measure Mechanism in [37]. Relatedly, [27] studied instance-optimal private density estimation under Wasserstein loss, and [38] extended these ideas to online synthetic data release.
Beyond Wasserstein distance, several alternative discrepancy measures have been considered, including $L_{1}$-type distances [61] and more general integral probability metrics (IPMs) [6]. The $1$-Wasserstein distance is a particular example of an IPM, and [12] characterized the sample complexity of DP synthetic data generation for unbounded query classes under IPMs. Moreover, [23] analyzed the class of $s$-sparse queries and obtained an $(\varepsilon n)^{-1/s}$ error rate.
@@
Can we generate DP synthetic data with utility guarantees for smooth queries that improve upon the $n^{-1/d}$ rate?
@@ 1.2 Main results
When $k=1$, since differentiable functions are dense in the class of 1-Lipschitz functions, this reduces to the $W_{1}$-distance between $p,q$ due to the Kantorovich-Rubinstein duality [57]. For any data set $X=(x_{1},\ldots,x_{n})\in\Omega^{n}$, where $\Omega=[-1,1]^{d}$, we denote its associated empirical distribution as
@@
Let $\varepsilon\in(0,1)$. There exists a polynomial-time $(\varepsilon,\delta)$-DP algorithm that outputs a synthetic dataset $Y\subseteq\Omega$ such that
$\mathbb{E}d_{k}(p_{X},p_{Y})\leq\begin{cases}C_{k}d\left(\dfrac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}{}\right)^{k/d}&\text{if }k<d,\\ \dfrac{C_{k}d\sqrt{\log(1.25/\delta)}\log(\varepsilon n)}{\varepsilon n}&\text{if }k=d,\\ \dfrac{C_{k}d\sqrt{\log(1.25/\delta)}}{\varepsilon n}&\text{if }k>d.\end{cases}$ (3)
@@
The detailed DP-synthetic data generation algorithm is presented in Algorithm 1. Theorem 1 uncovers a sharp phase transition in the role of smoothness: once the smoothness order exceeds the dimension ($k>d$), additional smoothness no longer yields further improvements in the error rate. In particular, for multivariate polynomial queries, including low-order moments and correlations, linear statistics, and quadratic functionals, the theorem establishes an error rate of $O_{k}(n^{-1})$, which is independent of the ambient dimension $d$. This strictly improves the $O\left(n^{-\frac{2k}{3d+2k}}\right)$ rate in [60, Theorem 22] when $k=1$, and recovers [48, Theorem 41], since differentiable functions are dense in the class of $1$-Lipschitz functions.
For simplicity, we assume $\varepsilon\in(0,1)$ in Theorem 1. When $\varepsilon>1$, the only modification is the choice of the variance in the Gaussian mechanism step of Algorithm 1. Based on Lemma 1, the same analysis applies, differing only in its dependence on $\varepsilon$ in the utility bound.
@@
The minimax lower bound is stronger and more general compared to the lower bound in [11], where the authors considered only $\varepsilon$-DP algorithms with the utility given by the 1-Wasserstein distance. When taking $\delta=0,k=1$, we recover the $n^{-1/d}$ lower bound in Wasserstein distance from [11]. Our proof is conceptually very different from the packing number argument in [11, Corollary 9.4].
@@
  • Vanishing $\delta$ (often $\delta=o(1/n)$). In fingerprinting-code and tracing-based lower bounds for private query answering (e.g., [16, 52]), the argument repeatedly applies approximate DP along a long adaptive process. As a result, the $\delta$-failure probability can accumulate (typically via union bounds) across many steps, and one therefore requires $\delta\ll 1/n$ to keep the overall failure probability negligible. A related requirement appears in the estimation lower-bound framework of [17], where a tracing attack combined with a group-privacy/union-bound step leads to the assumption $\delta=o(1/n)$.
  • $\varepsilon$-relative $\delta$ (no need for $\delta\to 0$). In DP-Le Cam and DP-Assouad [3], approximate DP enters primarily through an additive slack term that scales with the Hamming radius of the reduction (via group privacy), so it suffices to control $\delta$ relative to $\varepsilon$. For example, [3] yields a distinguishability contribution of the form $D\delta$; when the construction takes $D=\Theta(1/\varepsilon)$, assuming $\delta=O(\varepsilon)$ keeps this term bounded. Likewise, [42, Theorem 1.3] assumes $\delta\leq\varepsilon$ so that the accumulated approximate-DP slack remains $O(1)$.
@@
In the proof of Theorem 1, we leverage the $k$-smoothness of the function class $\mathcal{F}_{k}$ to achieve Chebyshev polynomial approximation rates superior to those provided by Jackson’s theorem for Lipschitz functions. In our setting, a generalized Jackson’s theorem for $k$-smooth functions is needed, and we provide an improved estimate based on [40] tailored to our setting in Lemma 3. In addition, we derive a higher-order global coefficient-decay result for multivariate Chebyshev expansions. Our primary technical contribution is an argument that translates these smoothness assumptions into global energy decay bounds (Theorem 3). While this result is central to our utility analysis of DP-synthetic data in Theorem 1, it may also be of independent interest for broader distribution approximation problems.
@@
The rest of the paper is organized as follows. We provide preliminary results on differential privacy and Chebyshev polynomials in Section 2. In Section 3, we establish a general theorem connecting the integral probability metric induced by $\mathcal{F}_{k}$ and Chebyshev moments between two probability measures. In Section 4, we present our DP-synthetic data generation algorithm and analyze its privacy and utility guarantees stated in Theorem 1. The proof of the minimax lower bound (Theorem 2) is given in Appendix A. Additional proofs are deferred to the Appendices B to E. The computational complexity of Algorithm 1 is analyzed in Appendix F.
@@ 3 Integral probability metric and Chebyshev moments
Next, we apply the result of the Chebyshev moments of $k$-smooth functions and study the relationship between two types of discrepancies of probability measures $p$ and $q$: the integral probability metric $d_{k}(p,q)$ and the weighted $\ell_{2}$-difference of their Chebyshev moments, stated as follows:
@@
Proof.
For any function $\in\mathcal{F}_{k}, ff$ is also $1$-Lipschitz. Then for the function $\tilde{f}=\sum_{K}\tilde{c}_{K}\overline{T}_{K}$, by Jackson’s Theorem introduced in Lemma 3, we have
@@ 4 DP synthetic data generation for $k$-smooth queries
3: Set $\sigma^{2}\geq\frac{4\cdot 2^{d}}{\pi^{d}}\cdot\frac{S\log(1.25/\delta)}{n^{2}\varepsilon^{2}}$, where $S=\sum_{K\in\{0,\dots,m\}^{d}}\frac{1}{\|K\|_{2}^{k}}$.
@@
Output: A synthetic dataset $Y\subseteq[-1,1]^{d}$ according to the probability measure $q=\sum_{J\in[r]^{d}}q_{J}\delta_{g_{J}}$, where $\delta$ is the Dirac delta function.
@@
    Compute the Chebyshev moments of the approximated distribution and add noise to guarantee privacy.
@@
    Generated a synthetic data distribution by matching the perturbed Chebyshev moments.
Note that in the output step of Algorithm 1, the size of the synthetic data $|Y|$ depends on the rational approximation of the density function of $q$, and we discuss the details here. Let $q_{J}$ be the weight of the probability measure $q$ on $g_{J}$. We can choose rational numbers $\{w_{J}\}_{J\in[r]^{d}}$ such that $\max_{J}|w_{J}-q_{J}|$ is arbitrarily small. Let $m^{\prime}$ be the least common multiple of the denominators of all $\{w_{J},J\in[r]^{d}\}$, then we output the synthetic dataset $Y$ containing $m^{\prime}w_{J}$ copies of $g_{J}$ for each $J$.
@@ 4.1 Privacy
where $f(X)_{K}$ denotes the $K$-th entry of the vector $f(X)_{K}$, and $\overline{T}_{K}({\mathbf{x}})$ is the $K$-th normalized multivariate Chebyshev polynomial.
Define $\operatorname{nnz}(K)$ to be the number of non-zero entries in $K$. Since $\max_{{\mathbf{x}}\in[-1,1]^{d}}|\overline{T}_{K}({\mathbf{x}})|\leq\sqrt{2^{\mathrm{nnz}(K)}/\pi^{d}}$, for adjacent datasets $X,X^{\prime}$, we have
@@
where $S=\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}$ as defined in Algorithm 1.
Let the $g(X)$ denote the rounded datasets computed Algorithm 1, i.e. $g(X)\coloneqq\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{n}\}$. Then for two adjacent datasets ${X},{X}^{\prime}$, we have $g(X),g(X^{\prime})$ are also adjacent. Thus we have
@@
By Gaussian mechanism (Lemma 1) and our choice of $\sigma^{2}$ in Algorithm 1, adding independent $N(0,\sigma^{2})$ to each component of $f\circ g(x)$ guarantees $(\varepsilon,\delta)$-differential privacy, which is exactly to output $(\|K\|_{2}^{-k/2}\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$. Then the moments $(\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$ satisfies $(\varepsilon,\delta)$-differential privacy.
Finally, since the remainder of Algorithm 1 simply post-processes $(\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$ without returning to the original data $X$, the output of the algorithm is also $(\varepsilon,\delta)$-differentially private.
@@ 4.2 Accuracy
Next, we show the accuracy bound of our synthetic data algorithm in the integral probability metric $d_{k}$. Let $\widetilde{{X}}$ denote the aggregated dataset $\{\tilde{x}_{1},\dots,\tilde{x}_{n}\}$ as the output. Then as each data point is moved by at most $\frac{\sqrt{d}}{2}\Delta$, there is
$d_{k}(\mu_{{X}},\mu_{\widetilde{{X}}})\leq\sup_{f:1\text{-Lip}}\big|\langle f,\mu_{{X}}-\mu_{\widetilde{{X}}}\rangle\big|=W_{1}(\mu_{{X}},\mu_{\widetilde{{X}}})\leq\frac{\sqrt{d}}{2}\Delta.$
Here $W_{1}$ denotes the 1-Wasserstein distance, and the equality is from (1). Since $q$ is optimal for construction under the weighted $\ell^{2}$-norm of the Chebyshev moments, there is
$\Gamma^{2}$ $\coloneqq\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 2\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}+\left|\widehat{m}_{K}-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 4\sum_{K}\frac{\left|\mathbb{E}_{X\sim\mu_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}}{\|K\|_{2}^{2k}}=4\sum_{K}\frac{\left|\eta_{K}\right|^{2}}{\|K\|_{2}^{2k}}.$
@@
$d_{k}(\mu_{\widetilde{{X}}},q)\leq\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\Gamma.$
@@
Applying the estimate of $S$ from Lemma 13 in Appendix E, we have
$\mathbb{E}\Gamma$ $\leq\frac{4\cdot 2^{d/2}}{\pi^{d/2}}\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot S$ (6)
@@
where $c>\sqrt{8/\pi}$ is an absolute constant. Therefore, we have an accuracy bound
$\mathbb{E}d_{k}(\mu_{{X}},q)$ $\;\leq\;\mathbb{E}d_{k}(\mu_{{X}},\mu_{\widetilde{{X}}})+\mathbb{E}d_{k}(\mu_{\widetilde{{X}}},q)\;\leq\;\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\mathbb{E}\Gamma.$
And in the three different cases, by choosing $m=\Big\lceil\frac{1}{c}\cdot\bigg(\frac{\varepsilon n}{\sqrt{\log(1.25/\delta)}}\bigg)^{1/\max\{d,k\}}\Big\rceil$, we have
@@
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot m^{d-k}$
@@
    $\mathbb{E}d_{k}(\mu_{{X}},q)$ $\leq\sqrt{d}\cdot\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log m$
    $\leq\;C_{k}^{\prime}d\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log(\varepsilon n);$

### v2

@@ 1 Introduction
Much of the existing DP literature focuses on designing task-specific mechanisms tailored to a predefined workload (e.g., a fixed set of queries or a particular model class). While often highly accurate, these approaches can require substantial expertise, ranging from sensitivity analysis and mechanism selection to careful privacy accounting across iterative computations. A complementary and increasingly practical alternative is differentially private synthetic data generation, where one releases a synthetic dataset that approximates statistical properties of the original data while satisfying differential privacy [32, 9, 61, 8]. By the post-processing property [26] of differential privacy, analysts can subsequently perform a wide range of downstream tasks on the released synthetic dataset without incurring additional privacy loss.
@@ 1.1 Differentially private synthetic data
Within this Wasserstein framework, [11] gave the first $\varepsilon$-DP algorithm achieving the minimax rate $n^{-1/d}$ (up to polylogarithmic factors) for data in the hypercube $0,1]^{d}$ under the $\ell_{\infty}$ metric. Subsequently, [37] developed two classes of algorithms that attain the same minimax-optimal $n^{-1/d}$ rate for $d\geq 2$. These constructions have enabled the analysis of several downstream learning tasks [31, 62, 18]. Under the relaxed notion of $(\varepsilon,\delta)$-DP, [48] analyzed a Chebyshev moment matching algorithm that achieves the same $n^{-1/d}$ Wasserstein rate.
Recent work has sought to surpass the worst-case $n^{-1/d}$ scaling by exploiting additional structure in the data. For example, [36] considered datasets concentrated near a $d^{\prime}$-dimensional subspace and obtained a rate $n^{-1/d^{\prime}}$ up to polynomial factors in $d$, while [23] studied data supported on an unknown $d^{\prime}$-dimensional manifold in $0,1]^{d}$. When the dataset satisfies suitable sparsity conditions, [39] showed that the error rate can be further improved via an adaptation of the Private Measure Mechanism introduced in [37]. In a related direction, [29] analyzed the practical Private Evolution algorithm [45] through its connection to the Private Signed Measure Mechanism in [37]. Relatedly, [27] studied instance-optimal private density estimation under Wasserstein loss, and [38] extended these ideas to online synthetic data release.
Beyond Wasserstein distance, several alternative discrepancy measures have been considered, including $L_{1}$-type distances [61] and more general integral probability metrics (IPMs) [6]. The $1$-Wasserstein distance is a particular example of an IPM, and [12] characterized the sample complexity of DP synthetic data generation for unbounded query classes under IPMs. Moreover, [23] analyzed the class of $s$-sparse queries and obtained an $n^{-1/s}$ error rate.
@@
Can we generate DP synthetic data with utility guarantees for smooth queries that improve the $n^{-1/d}$ rate?
@@ 1.2 Main results
When $k=1$, since $\mathcal{F}_{1}\cap\{1-\mathrm{Lipschitz~function}\}$ is dense in the class of 1-Lipschitz functions on $1,1]^{d}$ under the $\ell_{2}$-metric, this gives an upper bound on $W_{1}(p,q)$ due to the Kantorovich-Rubinstein duality [57].
For any data set $X=(x_{1},\ldots,x_{n})\in\Omega^{n}$, where $\Omega=[-1,1]^{d}$ equipped with the $\ell_{2}$-metric, we denote its associated empirical distribution as
@@
Let $\varepsilon\in(n^{-1},1)$. There exists a polynomial-time $(\varepsilon,\delta)$-DP algorithm that outputs a synthetic dataset $Y\subseteq\Omega$ such that
$\mathbb{E}d_{k}(p_{X},p_{Y})\leq\begin{cases}C_{k}d\left(\dfrac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}{}\right)^{k/d}&\text{if }k<d,\\ \dfrac{C_{k}\sqrt{\log(1.25/\delta)}\log(\varepsilon n)}{\varepsilon n}&\text{if }k=d,\\ \dfrac{C_{k}\sqrt{\log(1.25/\delta)}}{\varepsilon n}&\text{if }k>d.\end{cases}$ (3)
@@
The detailed DP-synthetic data generation algorithm is presented in Algorithm 1. Theorem 1 uncovers a sharp phase transition in the role of smoothness: once the smoothness order exceeds the dimension ($k>d$), additional smoothness no longer yields further improvements in the error rate. In particular, for multivariate polynomial queries, including low-order moments and correlations, linear statistics, and quadratic functionals, the theorem establishes an error rate of $O_{k}(n^{-1})$, which is independent of the ambient dimension $d$. This strictly improves the $O\left(n^{-\frac{2k}{3d+2k}}\right)$ rate in [60, Theorem 22] when $k=1$, and recovers [48, Theorem 41].
For simplicity, we assume $\varepsilon<1$ in Theorem 1. When $\varepsilon\geq 1$, the only modification is the choice of the variance in the Gaussian mechanism step of Algorithm 1. Based on Lemma 1, the same analysis applies, differing only in its dependence on $\varepsilon$ in the utility bound.
@@
Our minimax lower bound is more general than the lower bound in [11], which studies $\varepsilon$-DP algorithms with utility measured by the $W_{1}$ distance under the $\ell_{\infty}$ metric on $\Omega$. Since functions in $\mathcal{F}_{1}$ are $d$-Lipschitz with respect to the $\ell_{\infty}$ metric, specializing our result to $\delta=0$ and $k=1$ recovers the $n^{-1/d}$ Wasserstein lower bound of [11], up to a factor of $d$. Since obtaining a non-trivial bound $n^{-1/d}=O(1)$ requires $d=O(\log n)$, our result recovers [11] up to a $\log n$ factor. Finally, our proof strategy is fundamentally different from the packing-number argument used in [11, Corollary 9.4].
@@
  • Vanishing $\delta$. In fingerprinting-code and tracing-based lower bounds for private query answering (e.g., [16, 52]), the argument repeatedly applies approximate DP along a long adaptive process. As a result, the $\delta$-failure probability can accumulate (typically via union bounds) across many steps, and one therefore requires $\delta\ll 1/n$ to keep the overall failure probability negligible. A related requirement appears in the estimation lower-bound framework of [17], where a tracing attack combined with a group-privacy/union-bound step leads to the assumption $\delta=o(1/n)$.
  • $\varepsilon$-relative $\delta$. In DP-Le Cam and DP-Assouad [3], approximate DP enters primarily through an additive slack term that scales with the Hamming radius of the reduction (via group privacy), so it suffices to control $\delta$ relative to $\varepsilon$. For example, [3] yields a distinguishability contribution of the form $D\delta$; when the construction takes $D=\Theta(1/\varepsilon)$, assuming $\delta=O(\varepsilon)$ keeps this term bounded. Likewise, [42, Theorem 1.3] assumes $\delta\leq\varepsilon$ so that the accumulated approximate-DP slack remains $O(1)$.
@@
In the proof of Theorem 1, we leverage the $k$-smoothness of the function class $\mathcal{F}_{k}$ to achieve Chebyshev polynomial approximation rates superior to those provided by Jackson’s theorem for Lipschitz functions. In our setting, a generalized Jackson’s theorem for $k$-smooth functions is needed, and we provide an improved estimate based on [40] tailored to our setting in Lemma 3. In addition, we derive a higher-order global coefficient-decay result for multivariate Chebyshev expansions. Our primary technical contribution is an argument that translates these smoothness assumptions into global energy decay bounds (Theorem 3). While this result is central to our utility analysis of DP-synthetic data in Theorem 1, it may also be of independent interest for broader distribution approximation problems. Finally, in contrast to [48], which outputs a private probability measure, our algorithm produces an explicit differentially private synthetic dataset Y. This representation is more directly usable in downstream applications, and we explicitly quantify the size of Y in the proof of Theorem 1.
@@
The rest of the paper is organized as follows. We provide preliminary results on differential privacy and Chebyshev polynomials in Section 2. In Section 3, we establish a general theorem connecting the integral probability metric induced by $\mathcal{F}_{k}$ and Chebyshev moments between two probability measures. In Section 4, we present our DP-synthetic data generation algorithm and analyze its privacy and utility guarantees stated in Theorem 1. The proof of the minimax lower bound (Theorem 2) is given in Appendix A. Additional proofs are deferred to the Appendices B to F. The computational complexity of Algorithm 1 is analyzed in Appendix G.
@@ 3 Integral probability metric and Chebyshev moments
Next, we apply the result of the Chebyshev moments of $k$-smooth functions and study the relationship between two types of discrepancies of probability measures $p$ and $q$: the integral probability metric $d_{k}(p,q)$ and the weighted $\ell_{2}$-difference of their Chebyshev moments. The proof of Theorem 4 is deferred to Appendix E.
@@ 4 DP synthetic data generation for $k$-smooth queries
3: Set $\sigma^{2}\geq\frac{4\cdot 2^{d}}{\pi^{d}}\cdot\frac{S\log(1.25/\delta)}{n^{2}\varepsilon^{2}}$, where $S=\sum_{K\in\{0,\dots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}$.
@@
Define probability measure $q:=\sum_{J\in[r]^{d}}q_{J}\delta_{g_{J}}$.
5: Let $m^{\prime}=\lceil r^{d}\,(\varepsilon n)^{\min\{1,k/d\}}\rceil$, $a_{J}:=\lfloor m^{\prime}q_{J}\rfloor$ for $J\in[r]^{d}$ and $L:=m^{\prime}-\sum_{J\in[r]^{d}}a_{J}$. Increase $a_{J}$ by $1$ for the $L$ indices with the largest fractional parts of $m^{\prime}q_{J}$ (breaking ties arbitrarily).
Output: The multiset $Y$ of size $m^{\prime}$ containing exactly $a_{J}$ copies of $g_{J}$ for each $J\in[r]^{d}$.
@@
    Compute the Chebyshev moments and add noise to guarantee privacy.
@@
    Generated a synthetic data distribution $q$ by matching the perturbed Chebyshev moments.
  5. 5.
    Output a synthetic data set $Y$ where $p_{Y}$ approximates the distribution $q$.
@@ 4.1 Privacy
where $f(X)_{K}$ denotes the $K$-th entry of the vector $f(X)_{K}$, and $\overline{T}_{K}({\mathbf{x}})$ is the $K$-th normalized multivariate Chebyshev polynomial. Define $\operatorname{nnz}(K)$ to be the number of non-zero entries in $K$. Since $\max_{{\mathbf{x}}\in[-1,1]^{d}}|\overline{T}_{K}({\mathbf{x}})|\leq\sqrt{2^{\mathrm{nnz}(K)}/\pi^{d}}$, for adjacent datasets $X,X^{\prime}$, we have
@@
where $S=\sum_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}\frac{1}{\|K\|_{2}^{k}}$ as defined in Algorithm 1. Let the $g(X)$ denote the rounded datasets computed Algorithm 1, i.e. $g(X)\coloneqq\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{n}\}$. Then for two adjacent datasets ${X},{X}^{\prime}$, we have $g(X),g(X^{\prime})$ are also adjacent. Thus we have
@@
By Gaussian mechanism (Lemma 1) and our choice of $\sigma^{2}$ in Algorithm 1, adding independent $N(0,\sigma^{2})$ to each component of $f\circ g(x)$ guarantees $(\varepsilon,\delta)$-differential privacy, which is exactly to output $(\|K\|_{2}^{-k/2}\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$. Then the moments $(\widehat{m}_{K})_{K\in\{0,\ldots,m\}^{d}\setminus\{\mathbf{0}\}}$ satisfies $(\varepsilon,\delta)$-differential privacy. Finally, since the remainder of Algorithm 1 simply post-processes $(\widehat{m}_{K})$ without returning to the original data $X$, the output of the algorithm is also $(\varepsilon,\delta)$-differentially private.
@@ 4.2 Accuracy
Next, we show the accuracy bound under the integral probability metric $d_{k}$ in the following steps.
(1) Error in Step 2: Let $\widetilde{{X}}$ denote the aggregated dataset in Step 2. Since each data point is moved by at most $\frac{\sqrt{d}}{2}\Delta$ in $\ell_{2}$-distance, and $f\in\mathcal{F}_{k}$ is $\sqrt{d}$-Lipschitz under $\ell_{2}$-distance,
$d_{k}(p_{{X}},p_{\widetilde{{X}}})=\sup_{f\in\mathcal{F}_{k}}\big|\langle f,p_{{X}}-p_{\widetilde{{X}}}\rangle\big|\leq\sup_{i}\sup_{f\in\mathcal{F}_{k}}|f({\mathbf{x}}_{i})-f(\overline{{\mathbf{x}}}_{i})|\leq\sqrt{d}\sup_{i}\|{\mathbf{x}}_{i}-\overline{{\mathbf{x}}}_{i}\|\leq\frac{d\Delta}{2}.$
(2) Error in Step 4: Since $q$ is the optimal construction under the weighted $\ell^{2}$-norm of the Chebyshev moments,
$\Gamma^{2}$ $\coloneqq\sum_{K}\frac{\left|\mathbb{E}_{X\sim p_{\widetilde{{X}}}}\overline{T}_{K}(X)-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 2\sum_{K}\frac{\left|\mathbb{E}_{X\sim p_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}+\left|\widehat{m}_{K}-\mathbb{E}_{X\sim q}\overline{T}_{K}(X)\right|^{2}}{\|K\|_{2}^{2k}}$
$\leq 4\sum_{K}\frac{\left|\mathbb{E}_{X\sim p_{\widetilde{{X}}}}\overline{T}_{K}(X)-\widehat{m}_{K}\right|^{2}}{\|K\|_{2}^{2k}}=4\sum_{K}\frac{\left|\eta_{K}\right|^{2}}{\|K\|_{2}^{2k}}.$
@@
$d_{k}(p_{\widetilde{{X}}},q)\leq\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\Gamma.$
@@
Applying the estimate of $S$ from Lemma 13 in Appendix F, we have
@@
where $c>\sqrt{8/\pi}$ is an absolute constant.
(3) Error in Step 5: Recall
$m^{\prime}=\lceil r^{d}\,(\varepsilon n)^{\min\{1,k/d\}}\rceil=\lceil m^{kd}\,(\varepsilon n)^{\min\{1,k/d\}}\rceil$ (7)
and the definition of $a_{J}$ from Step 5 in Algorithm 1. Set $w_{J}:=\frac{a_{J}}{m^{\prime}}$ so that $\sum_{J\in[r]^{d}}w_{J}=1,\;w_{J}\geq 0$. Then we have
$\max_{J\in[r]^{d}}|w_{J}-q_{J}|\leq\frac{1}{m^{\prime}}\qquad\text{and}\qquad\sum_{J\in[r]^{d}}|w_{J}-q_{J}|\leq\frac{r^{d}}{m^{\prime}}.$ (8)
Let $p_{Y}=\sum_{J\in[r]^{d}}w_{J}\,\delta_{g_{J}}$, where $Y$ is the output dataset. Fix $f\in\mathcal{F}_{k}$. Since $\max_{1\leq i\leq d}\|\partial_{i}f\|_{\infty}\leq 1$, we have $\sup_{x,x^{\prime}\in[-1,1]^{d}}|f(x)-f(x^{\prime})|\leq\sum_{i=1}^{d}|x_{i}-x_{i}^{\prime}|\leq 2d$. Therefore, we may subtract a constant from $f$ (which does not change $\langle f,q-p_{Y}\rangle$) and assume $\|f\|_{\infty}\leq d$. Then
$\bigl|\langle f,q-p_{Y}\rangle\bigr|=\Big|\sum_{J\in[r]^{d}}(q_{J}-w_{J})\,f(g_{J})\Big|\leq\|f\|_{\infty}\sum_{J\in[r]^{d}}|q_{J}-w_{J}|\leq d\sum_{J\in[r]^{d}}|q_{J}-w_{J}|.$
Then from (8),
$d_{k}(q,p_{Y})\;\leq\;d\sum_{J\in[r]^{d}}|w_{J}-q_{J}|\;\leq\;d\,\frac{r^{d}}{m^{\prime}}\leq d(\varepsilon n)^{-\min\{1,k/d\}}.$
(4) Total error: Together, we have an accuracy bound
$\mathbb{E}d_{k}(p_{{X}},p_{Y})$ $\;\leq\;\mathbb{E}d_{k}(p_{{X}},p_{\widetilde{{X}}})+\mathbb{E}d_{k}(p_{\widetilde{{X}}},q)+\mathbb{E}d_{k}(q,p_{Y})$
$\leq\;d\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\mathbb{E}\Gamma+d(\varepsilon n)^{-\min\{1,k/d\}}.$
In the three different cases, by choosing $m=\Big\lceil\frac{1}{c}\cdot\bigg(\frac{\varepsilon n}{\sqrt{\log(1.25/\delta)}}\bigg)^{1/\max\{d,k\}}\Big\rceil$, we have
@@
    $\mathbb{E}d_{k}(p_{{X}},p_{Y})$ $\leq d\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot m^{d-k}+d(\varepsilon n)^{-k/d}$
@@
    $\mathbb{E}d_{k}(p_{{X}},p_{Y})$ $\leq d\Delta+\frac{2C_{k}^{\mathrm{Jac}}\cdot d}{m^{k}}+\sqrt{C_{k}}\cdot\frac{c^{d}\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log m+k(\varepsilon n)^{-1}$
    $\leq\;C_{k}^{\prime}\cdot\frac{\sqrt{\log(1.25/\delta)}}{\varepsilon n}\cdot\log(\varepsilon n);$